OpenStudy (anonymous):

How many grams of NH3 are needed to provide the same number of molecules as in 0.75 grams of SF6

OpenStudy (anonymous):

When you convert to number of moles, it accounts for how many molecules there are. NH3 has a molar mass of 17.03. SF6 has a molar mass of 146.06. The number of moles within 0.75 grams of SF6 is (0.75(g)/146.06(g/mol)) which equals 0.0051348 moles. 0.0051348*(NH3 molar mass) = 0.0051348 * 17.03 = .087445644 grams.